Please explain in detail, how to solve this please?
Solve the system of equations by substitution:
x + 2y = 9
3x - y = 13
Thank you for your assistance.

Answer by sarah_adam(201)

x + 2y = 9 --- eq 1
3x - y = 13 --- eq 2
3x+6y = 27 --- eq 1 * 3 ( multiply the entire equation with 3)
3x - y = 13 --- eq 2
(-) (+) (-)
-------------
7y = 14 --- (eq 1 - eq 2)
y = 14/7 = 2
substitute value of y in eq 1
x+2(2) = 9
x + 4 = 9
x = 9 - 4 = 5

by substitution method
x = 9 - 2y
substitute value of x in eq 2
3(9 - 2y) - y = 13
27 - 6y - y = 13
27 - 13 = 7y
14 = 7y
y = 14/7 =2
x = 9 - 2y
x = 9 - 2(2) = 9 - 4 = 5
x = 5 ; y = 2